# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

### v6

@@ Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational premise that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple application contexts. Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher, showing a significant collapse between average- to worst-case settings. We further demonstrate that correlations between sensor modalities reduce the worst-case entropy of using multiple sensors by up to $\approx$75compared to average-case Shannon entropy. This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits of min-entropy when combining 20 sensor modalities. These results raise the serious risk of attacks that exhaustively search the space of possible sensor measurements. Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security, and we strongly urge caution when relying on mobile sensor data for security applications.
@@ 1 Introduction
Modern mobile devices come equipped with an array of embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This rich sensor data has enabled applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] and reduce user authentication prompts [7, 8, 9, 10, 11].
@@
    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate, or brute force, the (joint) measurement space. This gives rise to fundamental security risks to schemes that rely on signals from single and multiple mobile sensors.
@@
    Ultimately, we advise against relying on commodity sensors as sources of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
@@ 3.1 Threat Model and Assumptions
We consider an adversary aiming to compromise sensor-based security schemes (e.g., key generation, proximity detection, and continuous authentication). The adversary may gather extensive statistical data about how smartphone sensors behave in everyday usage; for instance, from widely available open datasets. We assume the attacker focuses on predicting or guessing the sensor outputs by prioritising the most probable values first, exploiting any biases in the distribution of sensor measurements. As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough. We exclude capabilities such as fault injection and other hardware attacks (see [34]). While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as out-of-scope in this work. This threat model thus represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically. Our goal is to evaluate whether sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
@@ 3.2 High-level Methodology
    Relay [5]: Contains sensor measurements for approximately 1,500 NFC-based contactless transactions, each recorded at 100 Hz across several physical locations (e.g., cafés). The dataset encompasses accelerometer, gyroscope, and environmental readings taken in realistic payment scenarios.
@@ 4.3.2 Chow-Liu Approximation
Our framework evaluates the joint entropy over all sensor combinations. The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module. Processing all four datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
@@ 4.3.3 Results
The results imply that many sensors contribute redundant information, showing a fundamental limitation of multi-modal data in real-world devices. Combining signals increases the apparent capacity for unpredictability, but correlations between sensors means that the min-entropy from a large sensor ensemble is not be substantially higher than a reduced subset thereof. We note that standards use min-entropy as a safer, worst-case metric nowadays [26, 33]. It measures how close the distribution is to collapsing around the single most probable outcome that an adversary will target first. To the best of our knowledge, the entropy collapse brought about by highly correlated sensor modalities has not been before in existing work. We provide the full sensor combination results in our open-source repository.
@@ 5.1 Discussion
One possibility is to incorporate sensor-based randomness into an entropy pool that seeds a cryptographically secure pseudorandom number generator (CSPRNG). By doing so, a hybrid design can mitigate bias in the raw sensor data and produce outputs that pass statistical randomness tests (see the designs by Suciu et al. [27] and Wallace et al. [29]). However, such an approach implicitly contradicts the rationale for many existing sensor-based applications, which leverage similarity or correlation in sensor readings, such as matching motion traces for device pairing or zero-interaction authentication [42, 13, 2, 3, 4, 1]. Here, the desired property is not pure unpredictability but rather shared information between the sensor signals collected by one or more devices. We have shown that sensor signals inherently have low entropy, even when using multiple sensors simultaneously, exposing a fundamental security flaw in such designs.
We see no straightforward remedy for the entropy shortfall of mobile sensors. Extractors, user-assisted randomness, and entropy pools address different problems: reducing bias, injecting small amounts of fresh entropy, or improving output distribution. Nevertheless, they do not compensate for fundamentally weak signals, which we have shown to be the case across a range of modalities. Sensor data and its derivatives may still be appropriate for low-security purposes. However, relying on it alone for security applications is highly inadvisable due to the serious risk of adversaries who exploit statistical biases to reduce the effective search space.

### v7

@@ Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational premise that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple contexts. Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher, showing a significant collapse between average- to worst-case settings. We further demonstrate that correlations between sensor modalities reduce the worst-case entropy of using multiple sensors by up to $\approx$75compared to average-case Shannon entropy. This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits of min-entropy when combining 20 sensor modalities. Our results demonstrate that adversaries may feasibly predict sensor signals through an exhaustive, brute-force exploration of the entire measurement space. Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security, and we strongly urge caution when relying on mobile sensor data for security applications.
@@ 1 Introduction
Modern mobile devices are equipped with various embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This sensor data has enabled myriad applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] and reduce user authentication prompts [7, 8, 9, 10, 11].
@@
    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate, or brute force, the space of possible measurements. This gives rise to fundamental security risks to schemes that rely on signals from single and multiple mobile sensors.
@@
    Ultimately, we do not recommend relying on commodity sensors as a source of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
@@ 3.1 Threat Model and Assumptions
We consider an adversary aiming to compromise sensor-based security schemes (e.g., key generation, proximity detection, and continuous authentication). The adversary may gather extensive statistical data about how smartphone sensors behave in everyday usage; for instance, from widely available open datasets. We assume the attacker focuses on predicting or guessing the sensor outputs by prioritising the most probable values first, exploiting any biases in the distribution of sensor measurements. As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough. We exclude capabilities such as fault injection and other hardware attacks (see [34]). While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as outside the scope of this work. This threat model thus represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically. Our goal is to evaluate whether sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
@@ 3.2 High-level Methodology
    Relay [5]: Contains sensor measurements for approximately 1,500 NFC-based contactless transactions, each recorded at 100 Hz across several physical locations (e.g., cafés). The dataset contains accelerometer, gyroscope, and environmental readings from devices involved in emulated payment scenarios.
@@ 4.3.2 Chow-Liu Approximation
Our framework evaluates the joint entropy over all sensor combinations. The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module. Processing all datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
@@ 4.3.3 Results
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure 3: Entropy values versus the number of joint sensor modalities.
[…]
The results suggest that many sensors contribute redundant information when combined, exposing a fundamental limitation of multi-modal data in reality. Combining signals increases the apparent capacity for unpredictability, but correlations between sensors means that the min-entropy from a large sensor ensemble is not be substantially higher than a reduced subset thereof. We note that standards implore the use of min-entropy as a safer, worst-case metric [26, 33]. It measures how close the distribution is to collapsing around the single most probable outcome that an adversary will target first. To the best of our knowledge, the entropy collapse from highly correlated sensor modalities has not been elicited in existing work. We provide the full sensor combination results in our open-source repository.2
@@ 5.1 Discussion
One possibility is to incorporate sensor-based randomness into an entropy pool that seeds a cryptographically secure pseudo-random number generator (CSPRNG). By doing so, a hybrid design can mitigate bias in the raw sensor data and produce outputs that pass statistical randomness tests (see the designs by Suciu et al. [27] and Wallace et al. [29]). However, this approach implicitly contradicts the rationale for many existing sensor-based applications, which leverage similarity or correlation in sensor readings. For example, in matching similar motion traces or environmental conditions of two neighbouring devices [42, 13, 2, 3, 4, 1]. The desired property here is not pure unpredictability but rather shared information between the sensor signals collected by one or more devices. Using CSPRNG-derived signals would make it difficult to distinguish whether signals originated from co-located devices using similarity-based metrics. This exposes a fundamental security paradox in existing work. We have shown that sensors have insufficient entropy, even when using multiple ones simultaneously. Yet, methods to increase that entropy would compromise the usefulness of schemes that rely on the similarity of one or more sensor signals.
We see no straightforward remedy for the entropy shortfall of mobile sensors. Extractors, user-assisted randomness, and entropy pools address diHoweverfferent problems: reducing bias, injecting small amounts of fresh entropy, or improving output distribution. Nevertheless, they do not compensate for fundamentally weak signals, which we have shown to be the case across a range of modalities. Sensor data and its derivatives may still be appropriate for low-security purposes. However, relying on it alone for security applications is highly inadvisable due to the serious risk of adversaries who exploit statistical biases to reduce the effective search space.